OpenStudy (anonymous):

kendrick travels 2/3 mile to his friends house.He cycyles 3/4 of this distance and walks the rest of the way.What fraction of a mile did he cycle? can you break this down for me please step by step

OpenStudy (anonymous):

The distance between Kendrick's house and his friend's house is 2/3's of a mile. He cycles 3/4's of this way. You need to find the fraction of a mile that he cycled. If he cycled 3/4's of the 2/3's of the mile, then you just need to multiply the two fractions. \[\frac{ 3 }{ 4 } \times \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

6/12 = 1/2 ?

OpenStudy (anonymous):

Yeah, that's right!

OpenStudy (anonymous):

how did you know to multiply in that question? am not understanding how you figured that part out. why wouldn't it be adding or subtracting etc etc.

OpenStudy (anonymous):

thank you for your help btw. =)

OpenStudy (anonymous):

You have to look at the key words in the question. The first sentence says that the distance is 2/3's of a mile. That's one important number. In the next sentence, it says that he cycled 3/4's OF the distance. Whenever you see the word "of", that should clue you in that it is a multiplication problem. You also know the distance, 2/3 of a mile. So you can restate the question as "What is 3/4 of 2/3?" Because you know that "Of" is the same thing as multiplication, you can solve it.

OpenStudy (anonymous):

No problem :)